A current with sinusoidal time-dependence, whose peak value is 9.55 A and frequency is 359 cycles per second, passes through a long straight wire. A rectangular 18.2 cm
Answer #1

eavg = e = dBA/dt = A*(uo/2*pi*r)* dI / dt = (L1*L2*2*10^-7/0.202)*Imax*w = (0.182*0.0819*2*10^-7*9.55*359*2*3.14)/0.202 = 317.754e?6 v

P = e^2/2R = 1.071e?9/2 = 535.354e?12 W
Answer #2

H1 Mo 2tr

B_max = Uo * I_max / (2*pi*r) = 0.000009455445544554455 = 9.45545 micro tesla

B = B max * sin (2*pi*f*t) = 9.45545 sin ( 2255.66 t ) micro tesla

dB/dt = B_max* 2255.66 * cos (2255.66 t ) micro tesla /sec = 0.021328 cos (2255.66 t) tesla/sec

now

Area = 8.19*18.2*0.0001 = 0.01491 square meter

now emf induced = N*A*dB/dt = 0.01491*0.021328* cos (2255.66 t)

V = 0.0003179 cos (2255.66 t)

V_rms = V_max / sqrt(2) = 0.000224763 = 0.224763 milli volts

now average power = (V_rms)2/R = 5.3572 * 10^(-10) watts = 0.53572 nano watts